The following data was reported on total Fe for four types of iron formation (1 = carbonate, 2 = silicate, 3 = magnetite, 4 =The following data was reported on total Fe for four types of iron formation (1 = carbonate, 2 = silicate, 3 = magnetite, 4 = hematite).

1: 20.8 28.1 27.8 27.0 28.2 25.2 25.3 27.1 20.5 31.2
2: 26.9 24.0 26.2 20.2 23.5 34.0 17.1 26.8 23.7 24.6
3: 30.4 34.0 27.5 29.4 28.1 26.2 29.9 29.5 30.0 36.4
4: 36.4 44.2 34.1 30.3 31.8 33.1 34.1 32.9 36.3 25.8

Carry out an analysis of variance F test at significance level 0.01. State the appropriate hypotheses.

H0: μ1μ2μ3μ4
Ha: at least two μi's are equalH0: μ1 = μ2 = μ3 = μ4
Ha: at least two μi's are unequal    H0: μ1 = μ2 = μ3 = μ4
Ha: all four μi's are unequalH0: μ1μ2μ3μ4
Ha: all four μi's are equal


Summarize the results in an ANOVA table. (Round your answers two decimal places.)
Answer #1

Hypothesis ] Ho; - M, =l2 =M3= My Vis Hai; - all four is are unequal. Consider = carbonate Ig- magnetite - silicate Y4=hematiHere given whene ic treatments) {= 1, 2, 3,4 j=1,2,..., 10 Cobs?) Here Tai = 261.2 I Taj 26.12 i J j 247 24.7 Y. Σ 10 301.4 =S.S.Total = S.Stort t sis.E: sis.Ti = 517 7+ 563.6 = 108113 giren n=40 Wifi (Total) = n = 40-9=39 df (treatment) = K-1=4-1=3517 7 =172158 M.s.s.Trt = sisitot k-1 Mosis. Es S.S.E nok 563-3 15.66 36 14:56 = 11.02 Fratto - M.s.s.Tot M.S.S.E: ANONA TablDecision Relle; - Reject Ho If GIF a critical F Here in these case CONF > critica .. Reject wo Ho at a cool los. Conclusion4:22 44 白83% 1% significance level Fol(19) - 2 4 5 6 7 | 8 9 10 V2 2 3 5 6 9 10 11 12 13 14 15 16 17 18 19 20 21 22 24 26 27- Minitab - Untitled Eile Edit Data Calc Stat Graph Editor Tools Window Help Assistant TOHO % *SOMO 2 += WITO BO GF -- ASK 2- Minitab - Untitled Eile Edit Data Calc Stat Graph Editor Tools Window Help Assistant jij门的 BOGORE- %% 2 | +Y2 X Q TO Sessio